# Hidden long-range correlations in the ion distribution at the graphite / [bmim][NTf2] electrified interface

### v2

@@ Abstract
A capacitor consisting of the [bmim][NTf2] ionic liquid (IL) confined in between planar graphite electrodes has been investigated by molecular dynamics based on an all-atom, unpolarizable force field. Structural and dynamical properties such as: (i) the density and orientation of the [bmim]+ and [NTf2]- ions throughout the capacitor; (ii) the electrostatic double layer at the electrode / electrolyte interface; (iii) the ions’ mobility perpendicular and parallel to the graphite plates are determined as a function of the electrostatic charge of the capacitor, the concentration of absorbed water, the temperature and pressure. Grouping the [bmim]+ and [NTf2]- ions into neutral ion pairs reveals an intriguing ordering normal to the interface that is related to correlations among the dipole moments of the neutral ion pairs. These correlations might explain the observation of an anomalous Stark effect (Pockels effect) reported a few years ago in Langmuir 2021, 37, 5193-5201, and provides useful insight for the multitude of electro-chemical applications that involve electrode / ionic liquid interfaces.
@@ I Introduction
Since the strict confinement of the interfacial charge perturbation is supported by very robust arguments and is verified by a large number of experimental, theoretical and computational studies, we look for the responsible of this long range ordering at the second simpler and most likely suspect, represented by the of ions into neutral dipolar aggregates, which could display long range correlation effects. As the testbed for this analysis we have chosen the interface between a graphite electrode and a paradigmatic IL, consisting of the 1-butyl-3-methylimidazolium cation ([bmim]+) and bis(trifluoromethylsulfonyl)-imide anion ([NTf2]-), i.e., [bmim][NTf2]. This compound is mildly hydrophobic but retains a not-completely-negligible water concentration of up to about $25$ molar %, which however corresponds to only less that $1.5$ wgt% due to the mass disparity of water and [bmim][NTf2]. Samples of dry and water-contaminated [bmim][NTf2] samples with and without confining graphite electrodes have been simulated by molecular dynamics (MD) using an atomistic, semi-empirical force field model. We chose [bmim][NTf2] for our investigation because, despite its complexity, it has received a considerable interest as a solvent electrolyteelectrochem and because relatively abundant experimental and computational literature is available on its physical-chemistry properties. Moreover, the same effect might not be observable in simpler (either realistic or idealised) models because size and shape of the electrolyte ions might play a role in the propagation of correlation effects far from the interface. On the other hand, graphite has been selected because it is indeed used to make electrodes, and especially because its interaction with organic species are covered by the same type of interatomic force fields used for the ions themselves, unlike metal electrodes, whose typical empirical potential energy models are not easy to combines with those of ILs.
@@ II Model and Method
The subject of the major portion of the simulations is a planar capacitor consisting of two parallel graphite plates orthogonal to the $\hat{x}$ axis, delimiting a relatively thick ($\sim 11$ nm) [bmim][NTf2] layer, see Fig. 1. The thickness of the [bmim][NTf2] slab is such to overlap, at least to a limited extent, with the range (1010010-100 nm) of the long range correlations found in similar experimental systems. Both graphite and the IL are described at the atomistic level. The potential energy surface of the system is represented through an atomistic, empirical force field of the Amber form,amber parametrised at first using the Gromos model in its 54a754a7 version for the [bmim][NTf2] component, gromos and the SPC model for water.spc In practice, the topology files for the [bmim]+ and [NTf2]- ions have been obtained using the automated topology building utility made available by the ATB website.atb At the beginning, all the charges on the graphite carbon atoms are set to zero, to represent uncharged electrodes. In the case of electrodes of non-vanishing surface density $\sigma$, the atomic-like charges of graphite are distributed in contributions $\pm\delta q$ of equal size and opposite sign on the outermost layers of the graphite block, as detailed in the following Sec. III.2.
@@
Figure 3: [bmim][NTf2] neutral ion pair. The color coding is the same as in Fig. 1 The picture shows a generic configuration from a MD trajectory at $T=300$ K.
@@
The artificial periodicity along $x$, however, has a far more significant effect, that needs to be accounted for. An isolated, single planar capacitor of charge $Q=\sigma A$, where $\sigma$ is the surface charge density on the positive electrode and $A$ is the electrode area, is globally neutral. Hence, the electric field on its outside vanishes, and symmetry considerations show that inside the capacitor the electric field is oriented normal to the plates. Then, application of Gauss’ theorem immediately gives that the electric field strength on its inside is a constant equal to $4\pi\sigma$. For an infinitely periodic sequence of capacitors, however, the analysis is not so simple, because there is no simple way to locate a surface of vanishing electric field (${\bf E}=0$) from which to start applying Gauss’ theorem to compute ${\bf E}(x)$ anywhere else in space. Seen in a different but equivalent way, there is no reason why the electric field within the graphite blocks separating neighbouring replicas has to vanish. For these reasons, the electric field inside the liquid electrolyte, although still constant and directed along $x$, is not equal to $4\pi\sigma$.chris The precise value of ${\bf E}(x)$ can be computed analytically based on the Ewald sum formalism, but it depends on boundary condition parameters set in the simulation code, whose value in not easily determined. To avoid this ambiguity, the electric field $^{\bf E_{0}}=E_{0}(x)\hat{x}$ along the $x$ coordinate within the capacitor and within the graphite slab has been computed using Gromacs considering an empty capacitor whose charged plates (represented by the charge graphite layers) are located at the distance used in the actual simulations with the ionic liquid. This empty capacitor curve $E_{0}(x)$ has been used throughout the paper to compute the electrode contribution to the position-dependent electric field ${\bf E}(x)$ and electrostatic potential ${\bf\phi}(x)$ through the capacitor.
@@
All simulated systems, including the homogeneous liquid [bmim][NTf2] samples, consist of the same number of cations and anions. The electrolyte slab, therefore, is globally neutral. Since also the whole sample, which is 3D periodic, needs to be neutral, the two interfaces delimiting the electrolyte may only carry zero or opposite surface charge densities $\pm\sigma$. At first, the charge on each graphite C atom is identically zero. Then, any non-vanishing charge density on the graphite surfaces will be represented by atomic charges $q_{grf}=\pm\delta q$ of equal modulus and opposite sign on the two surfaces. Since graphite is a conductor, the static $\pm\sigma$ charges on the two sides of the graphite, in reality, would not be stable with respect to charge recombination. However, the electrical conductivity perpendicular to the graphite planes is about $10^{3}$ times lower than along the planes. Therefore, the simple charging model is justified by neglecting the electrical conductivity across the graphite slab, and assuming that the two outermost C-planes are connected to the two poles of a battery.
@@
Figure 4: Snapshot of the ionic liquid part of Fig. 1 coarse grained through Eq. 1. Red dots: cations; blue dots: anions. Temperature in measured in scaled energy units, see text.
@@
whose definition is apparently analogous to that of the centre of mass of a set of particles. The choice of emphasising the charges instead of the masses is due to our interest in the electrostatics of the sample. This definition is not as unanimously accepted nor extensively used as that of the centre of mass. In this context, however, it provides a simple way to coarse grain the charge distribution of [bmim][NTf2], which is fairly complex and difficult to visualise and interpret in its atomistic detail. For instance, the configuration of cations and anions corresponding to the snapshot in Fig. 1 is shown in Fig. 4, highlighting the drastic simplification of the IL structure achieved by the coarse graining expressed by Eq. 1.
The method to highlight subtle correlations in the distribution of ions in the simulated trajectories is somewhat complex and consists of a few steps. For this reason, the partition of ions into neutral pairs is defined in the following Sec. III using actual simulation results to illustrate this protocol.
@@ III.1 Properties of the homogeneous samples
The results show that the self-diffusion coefficient of cations and anions in pure, 3D homogeneous [bmim][NTf2] at $T=300$ K, although higher than that of the original Gromos 54a7 model, is still significantly below the experimental value measured by NMR.perk Moreover, and possibly more importantly, the estimated value depends on the time scale of the simulation, revealing a slightly sub-diffusive behaviour. Nevertheless, the mean square displacement is largely linear over a >50>50 ns interval, with an estimated value of $D$ of the order of $10^{-9}$ cm2/s for both anions and cations to within a large statistical error of $30$ %. This computed value has to be compared to the experimentally measured $10^{-7}$ $10^{-9}$ cm2/s, see the diffusion constant of [bmim][NTf2] measured by NMR as reported in Ref. perk, . We accept this discrepancy since unpolarizable ions are still the only way to simulate system sizes in excess of $10^{5}$ atoms at an acceptable cost. Addition of water at concentration as low as $1$ wgt% hardly affects the mobility of the ions. The self-diffusion coefficient of water in [bmim][NTf2] at 20 wgt% concentration and T=300T=300K turns out to be $(2.5\pm 0.5)\ 10^{-7}$ cm2/s, i.e., two orders of magnitude lower than in pure water at the same $T$, $P$ conditions.
@@ III.2 Structural, dynamic and electrostatic properties of the graphite / [bmim][NTf2] interface
The reference frame has been oriented in such a way that the interface is orthogonal to the $x$ axis. For this reason, local properties such as density and electrostatic potential profiles are plotted as a function of $x$. To simulate charged capacitors, small charges of up to a few $10^{-2}$ e (where e is the H+ charge) have been equally assigned to all the atoms of the two most external layers of graphite, qualitatively reflecting the picture provided by ab-initio (density functional theory) computations, see charging, . The charges are of equal magnitude and opposite sign on the two charged layers. Since also the IL slab is globally neutral, the whole sample is neutral. Since, according to the simulation results, the 2D number density of C atoms in the graphite layers at $T=300$ K is $37.08$ nm-2, a charge density of 0.05 e per graphite layer corresponds to a sizeable surface charge density of nearly $2$ e/nm-2.
@@
The charge profile across the capacitor is the required ingredient to compute the electrostatic potential as a function of $x$, displayed in Fig. 7 for the $T=300$ K sample at $\sigma=0$ and at $\sigma=2$ e/nm2. As expected, the electrostatic potentials is flat and the electric field vanishes over a wide central region of the capacitor, independently from the surface charge.
The output of the electrostatic potential analysis will provide data to compute the differential capacitance per unit area of the interface, defined as:
$C_{d}=\left(\frac{\partial\sigma}{\partial\varphi}\right)_{T,P}$ (2)
that represent a crucial parameter characterizing the electrostatic double layer at the interface, measuring the width of the double layer itself.
@@ III.3 Subtle electrostatic correlations deep into the bulk-like interior of the electrolyte
Acceptance - rejection of the bond interchange is decided by the usual Metropolis rule at fictitious temperature $T$, whose progressive reduction (simulated annealing) drives the system towards the minimum. The scale of he temperature interval and the schedule of the annealing have been decided by trial and error using a short sequence of preliminary computations. For the sake of simplicity, electrostatic and energy units will be expresses in scaled units, in which the unit of length is the nm, the unit of charge is $e$ and the unit of energy is $e^{2}/(1nm)$, corresponding to $0.0529$ Ha, or 1.44 eV, or $139$ kJ/mol. In these units, a unit dipole is $48$ Debye, a unit electric field is the one producing a unit energy change on a single positive $e$ charge upon a 1 nm displacement, etc. Also, measuring temperature in energy units, $T$ has been decreased from $T=2$ to $T=0.025$ energy units in 80 runs of $64\ 10^{6}$ MC steps, each consisting of the attempted interchange of bonds between neighbouring dipoles. Since the attempted MC move is a discrete event that cannot be fractioned, the acceptance ratio of the MC cannot be tuned, but it exceeded $0.5$ at high $T$, and decreased to $5\ 10^{-5}$ (one accepted move every 20 thousand moves) at the lowest $T=0.025$.
@@
The probability distribution for the separation $R_{d}$ of the charges for the dipoles in the final configuration is shown in Fig. 10. The average value $\langle R_{d}\rangle=0.5$ nm corresponds well to the geometric separation of cations and anions in the radial distribution function of the homogeneous [bmim][NTf2] liquid at the same conditions. The full width at half peak of about $0.1$ nm is sufficiently short to support the idea that the dipoles defined in this way are indeed physical associations of ions and not only fleeting geometrical details in the evolving ionic configuration.
@@
It can be supposed that a variety of subdivisions into pairs have comparable and relatively low self-energy, all representing the properties of the ideal unique and best pairing choice. However, repeated pairing optimisations, carried out by starting from different randomized starting point had a minimum fraction of equal pairs at the end of the simulated annealing of 66 %, showing that the self-energy basin around the (unknown) ideal minimum is relatively wide, but still well defined, without a plurality of significantly different local minima making ambiguous the pairing of ions into neutral pairs.
@@
To overcome this problem, the original ions’ population is supplemented by further ions, whose role is to represent the charge on the capacitor plates, providing a pairing opportunity for the counter-ions that populate the first IL layer next to the interface. This procedure is better illustrated with an example. Let us consider the same IL sample of $3072$ ion pairs, enclosed in between two graphite electrodes whose interfacial layers, each consisting of $4784$ carbon atoms, now carry $\pm 0.05\ e$ per C atom, for a total of $\pm 239.2$ e on the two capacitor plates. In the simulation, the left plate is negatively charged, and the right plate is positively charged. Since each carbon atoms carries only $\pm 0.05$ e, the representation of the charge on each electrode cannot be obtained by including all these atoms in the dipole formation protocol. Instead, a number of integer charges corresponding to the total charge on each electrode is introduced. In the present example $239$ e-e charges are located on the left electrode, and 239 +e+e charges are positioned on the right electrode.
At the beginning, cations are sorted in order of increasing distance from the anode. Then, each of the $239$ cations closest to the anode, of coordinates $\{{\bf r_{i}},i=1,...,239\}$ is paired with a e-e charge located $0.2$ nm below the graphite plane, at the same $(y_{i},z_{i})$ in-plane position of the corresponding cation. The specular operation is carried out with the opposite polarity electrode, interchanging the role of cations and anions, using the distance from the positively charged plate to select the anions to be paired with the surface positive charge.
@@
A quantitative assessment of the dipoles’ alignment along $x$ is given in Fig. 11, showing the average dipole moment along the directions parallel and normal to the interface. The data reported in this figure refer to which is displayed in Fig. 11 (b) for the sample illustrated in Fig. 13, having $\sigma=2$ surface charge density on the graphite electrodes. It is apparent that the distribution of the $d_{x}$ component of the dipole moment does not vanish over the entire with of the capacitor.

### v3

@@ Abstract
A capacitor consisting of the [bmim][NTf2] ionic liquid (IL) confined in between planar graphite electrodes has been investigated by molecular dynamics based on an all-atom, unpolarizable force field. Despite a few peculiarities due to the size and complexity of the ions, properties such as the density of ions throughout the capacitor, the screening of the surface charge on the electrodes by the IL and exact sum rules for the radial distribution functions of cations and anions generally comply with the results of time honored theories of the electrostatic double layer. This soothing regularity may conceal hidden correlations still compatible with the static screening rules, propagating far inside the IL the information on the state of charge of the capacitor. Evidence in this respect might have been detected by vibrational spectroscopy (see, for instance, Langmuir 2021, 37, 5193-5201) showing changes in optical properties of the IL far from the charged electrodes. We show that grouping the [bmim]+ and [NTf2]- ions into instantaneous neutral pairs reveals an intriguing long range ordering of ions normal to the interface, driven by the capacitor state of charge. These correlations manifest themselves through the parallel orientation of the dipole moments of the neutral ion pairs. We speculate that this effect changes the average intensity of fluctuating electric fields deep in the IL, while average, static fields vanish in agreement with well established screening laws. This effect, which could change the spectroscopic properties of the IL, is present in the simulated [bmim][NTf2] / graphite capacitor, but too small to be unambiguously confirmed by the present simulations with a safe margin over the error bar. The conceptual interest in these effects, however, will motivate further studies of the same or similar electrode / ionic liquid interfaces.
@@ I Introduction
Figure 2: Atomistic structure of the: (a) [bmim]+ and: (b) [NTf2]- ions. The color coding is the same as in Fig. 1 The picture shows a generic configuration from a MD trajectory at $T=300$ K. The gray circle with the plus and minus sign are located at the position of the coarse grained representation of the ions obtained by applying Eq. 1.
@@
Since the strict confinement of the interfacial charge perturbation is supported by very robust arguments and is verified by a large number of experimental, theoretical and computational studies, we look for the responsible of this long range ordering at the second simpler and most likely suspect, represented by the combination of ions into neutral dipolar aggregates, which might display long range correlation effects. As the testbed for this analysis we have chosen the interface between a graphite electrode and a paradigmatic IL, consisting of the 1-butyl-3-methylimidazolium cation ([bmim]+) and bis(trifluoromethylsulfonyl)-imide anion ([NTf2]-), i.e., [bmim][NTf2]. This compound is mildly hydrophobic but retains a not-completely-negligible water concentration of up to about $25$ molar %, which however corresponds to only less that $1.5$ wgt% due to the mass disparity of water and [bmim][NTf2]. Samples of dry and water-contaminated [bmim][NTf2] samples with and without confining graphite electrodes have been simulated by molecular dynamics (MD) using an atomistic, semi-empirical force field model. We chose [bmim][NTf2] for our investigation, despite its complexity, because it has received a considerable interest as a solvent electrolyteelectrochem and because relatively abundant experimental and computational literature is available on its physical-chemistry properties. Moreover, the same effect might not be observable in simpler (either realistic or idealised) models because size and shape of the electrolyte ions might play a role in the propagation of correlation effects far from the interface. On the other hand, graphite has been selected because it is indeed used to make electrodes, and especially because its interaction with organic species are covered by the same type of interatomic force fields used for the ions themselves, unlike metal electrodes, whose typical empirical potential energy models are not easy to combines with those of ILs.
@@ II Model and Method
The subject of the major portion of the simulations is a planar capacitor consisting of two parallel graphite plates orthogonal to the $\hat{x}$ axis, delimiting a relatively thick ($\sim 11$ nm) [bmim][NTf2] layer, see Fig. 1. The thickness of the [bmim][NTf2] slab is such to overlap, at least to a limited extent, with the range (1010010-100 nm) of the long range correlations found in similar experimental systems. Both graphite and the IL are described at the atomistic level. The potential energy surface of the system is represented through an atomistic, empirical force field of the Amber form,amber parametrised at first using the Gromos model in its 54a754a7 version for the [bmim][NTf2] component, gromos and the SPC model for water.spc In practice, the topology files for the [bmim]+ and [NTf2]- ions, whose structure is shown in Fig. 2, have been obtained using the automated topology building utility made available by the ATB website.atb At the beginning, all the charges on the graphite carbon atoms are set to zero, to represent uncharged electrodes. In the case of electrodes of non-vanishing surface density $\sigma$, the atomic-like charges of graphite are distributed in contributions $\pm\delta q$ of equal size and opposite sign on the outermost layers of the graphite block, as detailed in the following Sec. III.2.
@@
The artificial periodicity along $x$, however, has a far more significant effect, that needs to be accounted for. An isolated, single planar capacitor of charge $Q=\sigma A$, where $\sigma$ is the surface charge density on the positive electrode and $A$ is the electrode area, is globally neutral. Hence, the electric field on its outside vanishes, and symmetry considerations show that inside the capacitor the electric field is oriented normal to the plates. Then, application of Gauss’ theorem immediately gives that the electric field strength on its inside is a constant equal to $4\pi\sigma$. For an infinitely periodic sequence of capacitors, however, the analysis is not so simple, because there is no simple way to locate a surface of vanishing electric field (${\bf E}=0$) from which to start applying Gauss’ theorem to compute ${\bf E}(x)$ anywhere else in space. Seen in a different but equivalent way, there is no reason why the electric field within the graphite blocks separating neighbouring replicas has to vanish. Hence, the electric field inside the liquid electrolyte, although still constant and directed along $x$, is not equal to $4\pi\sigma$.chris The precise value of ${\bf E}(x)$ can be computed analytically based on the Ewald sum formalism, but it depends on boundary condition parameters set in the simulation code, whose value in not easily determined. To avoid this ambiguity, the electric field $^{\bf E_{0}}=E_{0}(x)\hat{x}$ along the $x$ coordinate within the capacitor and within the graphite slab has been computed using Gromacs considering an empty capacitor whose charged plates (represented by the charge graphite layers) are located at the distance used in the actual simulations with the ionic liquid. This empty capacitor curve $E_{0}(x)$ has been used throughout the paper to compute the electrode contribution to the position-dependent electric field ${\bf E}(x)$ and electrostatic potential ${\bf\phi}(x)$ through the capacitor.
@@
All simulated systems, including the homogeneous liquid [bmim][NTf2] samples, consist of the same number of cations and anions. The electrolyte slab, therefore, is globally neutral. Since also the whole sample, which is 3D periodic, needs to be neutral, the two interfaces delimiting the electrolyte may only carry zero or opposite surface charge densities $\pm\sigma$. At first, the charge on each graphite C atom is identically zero. Then, any non-vanishing charge density on the graphite surfaces will be represented by atomic charges $q_{gr}=\pm\delta q$ of equal modulus and opposite sign on the two surfaces. Since graphite is a conductor, the static $\pm\sigma$ charges on the two sides of the graphite, in reality, would not be stable with respect to charge recombination. However, the electrical conductivity perpendicular to the graphite planes is about $10^{3}$ times lower than along the planes. Therefore, the simple charging model is justified by neglecting the electrical conductivity across the graphite slab, and assuming that the two outermost C-planes are connected to the two poles of a battery.
@@
Figure 4: (a) Snapshot of the ionic liquid part of Fig. 1 coarse grained through Eq. 1. Red dots: cations; blue dots: anions. (b) radial distribution functions of cations and anions at the same conditions of panel (a). Only the ions in the central $5$ nm of the ionic liquid slab have been considered as centres to compute the corresponding $g_{\alpha\beta}(r)$.
@@
whose definition is apparently analogous to that of the centre of mass of a set of particles. The choice of emphasising the charges instead of the masses is due to our aim of identifying each ion with its charge. This definition is not as unanimously accepted nor extensively used as that of the centre of mass. In this context, however, it provides a simple way to coarse grain the charge distribution of [bmim][NTf2], which is fairly complex and difficult to visualise and interpret in its atomistic detail. For instance, the configuration of cations and anions corresponding to the snapshot in Fig. 1 is shown in Fig. 4 (a), highlighting the drastic simplification of the IL structure achieved by the coarse graining expressed by Eq. 1. The radial distribution functions (rdf) computed for the binary mixture consisting of cations and anions and shown in Fig. 4 (b) present all the features expected for a Coulombic fluid, including: the charge alternation around each ion chosen as the origin of the rdf; a tall anion-cation $g_{+-}$ peak, and a strong non-additivity in the distance of closest approach for the ++++ (long), ++- (short) and -- (long) ion pairs. The coordination numbers ($\sim 6$ for cation-anion and $\sim 12$ for anion-anion, while the cation-cation cannot be distinguished) are vaguely reminiscent of the NaCl structure.
The method to highlight subtle correlations in the distribution of ions in the simulated trajectories is somewhat complex and consists of a few successive steps. For this reason, the partition of ions into neutral pairs is defined in the following Sec. III using actual simulation results to illustrate this protocol.
@@ III.1 Properties of the homogeneous samples
The results show that the self-diffusion coefficient of cations and anions in pure, 3D homogeneous [bmim][NTf2] at $T=300$ K, although higher than that of the original Gromos 54a7 model, is still significantly below the experimental value measured by NMR.perk Moreover, and possibly more importantly, the estimated value depends on the time scale of the simulation, revealing a slightly sub-diffusive behaviour. Nevertheless, the mean square displacement is largely linear over a >50>50 ns interval, with an estimated value of $D$ of the order of $10^{-9}$ cm2/s for both anions and cations to within a large statistical error of $30$ %. This computed value has to be compared to the experimentally measured $10^{-7}$ cm2/s, see the diffusion constant of [bmim][NTf2] measured by NMR as reported in Ref. perk, . We accept this discrepancy since unpolarizable ions are still the only way to simulate system sizes in excess of $10^{5}$ atoms at an acceptable cost. Addition of water at concentration as low as $1$ wgt% hardly affects the mobility of the ions. The self-diffusion coefficient of water in [bmim][NTf2] at 20 wgt% concentration and T=300T=300K turns out to be $(2.5\pm 0.5)\ 10^{-7}$ cm2/s, i.e., two orders of magnitude lower than in pure water at the same $T$, $P$ conditions.
@@ III.2 Structural, dynamic and electrostatic properties of the graphite / [bmim][NTf2] interface
The reference frame has been oriented in such a way that the interface is orthogonal to the $x$ axis. For this reason, local properties such as density and electrostatic potential profiles are plotted as a function of $x$. To simulate charged capacitors, small charges of up to a few $10^{-2}$ e (where e is the H+ charge) have been equally assigned to all the atoms of the two most external layers of graphite, qualitatively reflecting the picture provided by ab-initio (density functional theory) computations, see Ref. charging, . The charges are of equal magnitude and opposite sign on the two charged layers. Since also the IL slab is globally neutral, the whole sample is neutral. Since, according to the simulation results, the 2D number density of C atoms in the graphite layers at $T=300$ K is $37.08$ nm-2, a charge density of 0.05 e per graphite layer corresponds to a sizeable surface charge density of nearly $2$ e/nm-2.
@@
The simukation results for the number and charge density profile as a function of $x$ across the capacitor are shown in Fig. 5 and Fig. 6. The charge profile across the capacitor, in particular, is the required ingredient to compute the electrostatic potential as a function of $x$, displayed in Fig. 7 for the $T=300$ K sample at $\sigma=0$ and at $\sigma=2$ e/nm2. As expected, the electrostatic potentials is flat and the electric field vanishes over a wide central region of the capacitor, independently from the surface charge.
@@ III.3 Subtle electrostatic correlations deep into the bulk-like interior of the electrolyte
Acceptance - rejection of the bond interchange is decided by the usual Metropolis rule at fictitious temperature $T$, whose progressive reduction (simulated annealing) drives the system towards the minimum. The scale of he temperature interval and the schedule of the annealing have been decided by trial and error using a short sequence of preliminary computations. For the sake of simplicity, electrostatic and energy units will be expresses in scaled units, in which the unit of length is the nm, the unit of charge is $e$ and the unit of energy is $e^{2}/(1nm)$, corresponding to $0.0529$ Ha, or 1.44 eV, or $139$ kJ/mol. In these units, a unit dipole is $48$ Debye, a unit electric field is the one producing a unit energy change on a single positive $e$ charge upon a 1 nm displacement, etc. Also, measuring temperature in energy units, $T$ has been decreased from $T=2$ to $T=0.025$ energy units in 80 runs of $64\ 10^{6}$ MC steps, each consisting of the attempted interchange of bonds between neighbouring dipoles. Since the attempted MC move is a discrete event that cannot be fractioned, the acceptance ratio of the MC cannot be tuned, but it exceeded $0.5$ at high $T$, and decreased to $5\ 10^{-5}$ (one accepted move every 20 thousand moves) at the lowest temperature, i.e., $T=0.025$.
@@
The probability distribution for the separation $R_{d}$ of the charges for the dipoles in the final configuration is shown in Fig. 10. The average value $\langle R_{d}\rangle=0.5$ nm corresponds well to the geometric separation of cations and anions in the radial distribution function of the homogeneous [bmim][NTf2] liquid at the same conditions. The full width at half peak of about $0.1$ nm is sufficiently short to support the idea that the dipoles defined in this way, although not permanent, are indeed physical associations of ions and not only fleeting geometrical details in the evolving ionic configuration of the [bmim][NTf2] sample.
It can be supposed that a variety of subdivisions into pairs have comparable and relatively low self-energy, all representing the properties of the ideal unique and best pairing choice. However, repeated pairing optimisations, carried out by starting from different randomized starting points had a minimum fraction of equal pairs at the end of the simulated annealing of 66 %, showing that the self-energy basin around the (unknown) ideal minimum is relatively wide, but still well defined, without a plurality of significantly different local minima making too ambiguous the pairing of ions into neutral pairs.
@@
To overcome this problem, the original ions’ population is supplemented by further ions, whose role is to represent the charge on the capacitor plates, providing a pairing opportunity for the counter-ions that populate the first IL layer next to the interface. This procedure is better illustrated with an example. Let us consider the same IL sample of $3072$ ion pairs, enclosed in between two graphite electrodes whose interfacial layers, each consisting of $4784$ carbon atoms, now carry $\pm 0.05\ e$ per C atom, for a total of $\pm 239.2$ e on the two capacitor plates. In the simulation, the left plate is negatively charged, and the right plate is positively charged. Since each carbon atoms carries only $\pm 0.05$ e, the representation of the charge on each electrode cannot be obtained by including all these atoms in the dipole formation protocol. Instead, a number of integer charges corresponding to the total charge on each electrode is introduced. In the present example $239$ (e)(-e) charges are located on the left electrode, and 239 $(+e)$ charges are positioned on the right electrode.
At the beginning, cations are sorted in order of increasing distance from the anode. Then, each of the $239$ cations closest to the anode, of coordinates $\{{\bf r_{i}},i=1,...,239\}$ is paired with a (e)(-e) charge located $0.2$ nm below the graphite plane, at the same $(y_{i},z_{i})$ in-plane position of the corresponding cation. The specular operation is carried out with the opposite polarity electrode, interchanging the role of cations and anions, using the distance from the positively charged plate to select the anions to be paired with the surface positive charge.
@@
A quantitative assessment of the dipoles’ alignment along $\hat{x}$ is given in Fig. 12, showing the average dipole moment density along the directions parallel and normal to the interface. The data reported in this figure refer to the dipole configurations shown in Fig. 13. Panel (a) in Fig. 12, in particular, refers to the $\sigma=0$ sample in Fig. 13, while Panel (b) in Fig. 12 refers to the $\sigma=2$ $e/nm^{2}$ sample in the same Fig. 13. The $\langle d_{y}\rangle$ and $\langle d_{z}\rangle$ components of the dipoles, on average, vanish across the whole capacitor in both samples. Again in both samples, the $\langle d_{x}\rangle$ component is affected by the presence of the interface over a short contact layer of sub-nanometric width. The important novelty is that $\langle d_{x}\rangle$ does not vanish well inside the [bmim][NTf2] layer in the $\sigma=2$ $e/nm^{2}$ case.
III.4 Speculations on the mechanism connecting the dipoles’ configuration and the spectroscopic properties
Figure 14: Probability distribution for the Cartesian components of the force $\bf F_{\alpha}$} acting on each ion pair $\alpha$ in the cental portion of the [bmim][NTf2] slab (see text) due to all other ion pairs in the sample. Comparison of the results for the capacitor at: $\sigma=0$ and $\sigma=2$ $e/nm^{2}$ surface charge density.
The discovery of a geometric parameter such as $\langle d_{x}\rangle$ whose non-vanishing value carries information on the presence of a charged interface well inside the [bmim][NTf2] slab represents a novel and interesting observation, which goes beyond the established description of electrified interfaces. In itself, however, this observation does not explain the changes in spectroscopic properties revealed by a few experiments. In fact, the computational and experimental results might have no obvious connection with each other apart from probing the same system. Nevertheless, it is conceivable that the parallel orientation of instantaneous dipoles due to a charge on the capacitor could reveal a deeper aspect which could explain both the dipoles ordering and the change in optical properties. This linking aspect needs to be compatible with the classical results of static screening, but otherwise could affect the ionic liquid properties as far as the correlation of the instantaneous dipoles could penetrate. A conceivable scenario concerns the properties of the fluctuating electric fields acting on the ions and evolving according to their thermal motion. Far from the graphite electrodes, the average electric field has to vanish, since [bmim][NTf2] is a conductor, but no obvious constraint limits the size as well as the position and size variations of fluctuating electric fields.
To verify this scenario, we identified all dipolar IL ion pairs whose cation is (instantaneously) located within the central $5$ nm of the capacitor. For each pair, the total electrostatic force acting on the cation and corresponding anion was computed, accounting for all other ions in the sample, i.e., not only those in the central $5$ nm slab. The surface charge density on the electrodes is not included in this computation, whose target is the quantification of the fluctuating electric fields in the bulk region of the ionic liquid. Since the simulated samples are large, the electrostatic force on the ions accounted for the complementary error part of the Ewald energy only, neglecting all reciprocal space terms. This approximation is introduced for the sake of simplicity, but its accuracy is enhanced by selecting the decaying length of the complementary error function in such a way to prevent the interaction of the central ions with the interfaces, but otherwise to be as large as possible to minimize the size of reciprocal space terms.
The computation of the electric fields has been carried out on a population of independent configurations for the samples at $\sigma=0$ and $\sigma=2$ $e/nm^{2}$. The results are illustrated in Fig. 14. They can be summarised as follows. In both cases, the average force vanishes to the same accuracy in the three Cartesian directions, confirming that the screening of external electric fields is verified to within the accuracy of the computation. The properties of electric field fluctuating in time and space are characterised by computing the distribution probability for the intensity of the three Cartesian components of the force on each dipole. Both for the $\sigma=0$ and the $\sigma=2$ $e/nm^{2}$ capacitors, the probability distribution for the forces along $\hat{y}$ and $\hat{z}$ are the same, again to within the computational accuracy, and the values of the two probability distributions are averaged together to improve the statistics in the comparison with the result in the $\hat{x}$ directions. As apparent in Fig. 14, the probability distribution for $|F_{x}|$ cannot be distinguished from the $(|F_{y}|+|F_{z}|)/2$ case for the uncharged capacitor. A slight difference appears in the charged capacitor case, localised in the range of low fields. The height of the two peaks differ by about 88%, which is only slightly less than their combined error bar. As a consequence of the slight difference in $P(F)$, the average value of $|F_{x}|$ is about $1.5$ % lower than the average value of $(|F_{y}|+|F_{z}|)/2$, again comparable to the combined error bar on the two quantities that are being compared. Increasing statistics to decrease the statistical error bar well beyond the separation of these quantities would require carrying out simulations and configuration analyses at least for times longer than those already done, a target that would require a dedicated effort as well as dedicated computational resources. To be safe, we can state that, at this stage, the computation is not conclusive, neither definitely supporting and certainly not disproving that the geometric properties of the dipoles population and the spectroscopic properties are related by changes in the fluctuating electric field inside the ionic liquid. Nevertheless, the scenario outlined in this section remains appealing, since it does not violate any known condition and it is not disproved by the present simulations.